1. In 2002 in the U.S. the public debt was $6.2 trillion. Put this number in perspective by relating the debt to GDP, to other countries’ debt, to the amount of interest payments on the debt, and to ownership of the debt.
Answer #1

IN 2002 US PUBLIC DEBT WAS $6.2 TRILLION AND GDP OF US ECONOMY WAS 10.928 LAC CRORES THAT IS EQUAL TO $ 10.928 TRILLION.

THEREFORE DEBT TO GDP RATIO COMES TO $6.2/10.928 = 56.47%

THE OTHER COUNTRIES DEBT AS OF 2002 TERMED AS PUBLIC DEBT WAS $3.5 TRILLION

THUS US DEBT COMPARED TO PUBLIC DEBT IS $6.2/3.5 IS 1.77 TIMES

WHILE COMPARED TO GDP AS A PERCENTAGE IS 3.5/10.928= 32.04%

INTEREST PAID DURING THE YEAR 2002 WAS 332.50 BILLION DOLLARS OR 0.333 TRILLION DOLLARS THUS IN THIS PERSPECTIVE COMPARED TO US DEBT THE PERCENTAGE OF INTEREST IS 0.333/6.2*100 = 5.34%

OWNERSHIP OF DEBT :- LEADING HOLDERS OF US DEBT ARE JAPAN AND CHINA . AS OF 2002 CHINA HELD 1.113 TRILLION DOLLAR OUT OF 6.2 TRILLION DOLLARS WHILE JAPAN HELD 1.064 TRILLION DOLLARS. THAT AMOUNTS TO 17% AND 16.5 %

LEADING HOLDERS OF US DEBT
COUNTRY DEBT HELD PERCENTAGE
IN TRILLION DOLLARS
CHINA                                   1.11                                 17.30
JAPAN                                   1.06                                 16.54
BRAZIL                                   0.31                                    4.77
UK                                   0.30                                    4.66
IRELAND                                   0.27                                    4.20
SWITZERLAND                                   0.23                                    3.53
LUXEMBOURG                                   0.22                                    3.48
CAYMAN ISLANDS                                   0.22                                    3.37
HONG KONG                                   0.20                                    3.11
BELGIUM                                   0.18                                    2.80
SAUDI ARABIA                                   0.18                                    2.75
TAIWAN                                   0.17                                    2.66
INDIA                                   0.16                                    2.41
SINGAPORE                                   0.14                                    2.16
FRANCE                                   0.13                                    2.02
SOUTH KOREA                                   0.12                                    1.79
CANADA                                   0.10                                    1.59
NORWAY                                   0.10                                    1.51
OTHERS                                   1.25                                 19.35
TOTAL                                   6.43                               100.00